OpenStudy (anonymous):

PLS HELP will give medal Arrange these numbers from least to greatest: 5/6 ,7/9, 3/5 A.5/6, 7/9 ,3/5 B.5/6 ,7/9 ,3/5 C.3/5,7/9,5/6 D.3/5,5/6,7/9

OpenStudy (anonymous):

@Haseeb96

OpenStudy (anonymous):

is it D

OpenStudy (anonymous):

you could multiply all the numbers with the MCM of 6-9-5

OpenStudy (haseeb96):

just take their values in decimals and then arrange them from least to greater

OpenStudy (haseeb96):

*take out

OpenStudy (anonymous):

oo ok

OpenStudy (anonymous):

so its c

OpenStudy (anonymous):

(5/6)*90=150 (7/9)*90=70 (3/5)*90=54 is more easy to arrange the numbers 54<70<150

OpenStudy (anonymous):

so * them all by 90 can you do that in other questions

OpenStudy (anonymous):

no, Your multiply to 90 because there is the MCM of 6,9,5